# Vintix: Action Model via In-Context Reinforcement Learning

### v1

@@ 1 Introduction
In this work, we start an investigation into cross-domain action models through In-Context Reinforcement Learning (ICRL), grounded in the Algorithm Distillation framework (Laskin et al., 2022). By focusing on a data-centric approach to ICRL (summarized in Figure 2), we present Vintix, a multi-task action model that exhibits initial signs of inference-time self-correction on training tasks (Figure 1) and preliminary evidence of adaptation to parametric task variations (Section 3.3).
@@
Continuous Noise Distillation (Section 2.1): We propose a data collection strategy extending the work of Zisman et al. (2024a) to continuous action spaces to ease the collection of training data.
Open Tools and Datasets for ICRL (Section 2.2): We publicly release datasets for 87 tasks across four domains (Meta-World, MuJoCo, Bi-DexHands, Industrial-Benchmark), along with data collection tools and instrumentation to support the development of action models eliciting ICRL behavior.
@@ 2 Approach
At the core of our approach (Figure 2) is Algorithm Distillation (Laskin et al., 2022), a two-step Offline Meta-RL algorithm. The first step involves collecting ordered training histories from base reinforcement learning (RL) algorithms, while the second step involves a decoder-only transformer trained solely for the next-action prediction. This approach facilitates in-context learning by effectively distilling the policy improvement operator into a causal sequence model. We further propose two augmentations to this technique: (1) democratizing the data collection process by introducing a continuous extension of the noise-distillation procedure by Zisman et al. (2024a); (2) conducting generalist agent-style cross-domain training on the acquired dataset.
@@ 2.1 Continuous Noise Distillation
The data collection strategy was originally designed for discrete action spaces. In our work, we propose its extension to continuous action spaces, where the resulting action is defined as a linear mixture of uniform random noise and demonstrator actions. This contrasts with Brown et al. (2019), who employed an epsilon-greedy approach that alternates between fully random and fully expert actions with probability . The pseudo-code for our noise-distillation procedure is formalized in Algorithm 1.
The noise schedule plays a crucial role in the generation of training trajectories. We observed that linear annealing of often results in non-smooth trajectories with abrupt transitions from random to demonstrator performance. This effect can negatively impact the model’s convergence; therefore, careful tuning of the noise schedule was necessary for certain tasks. A more detailed description of the epsilon decay functions is provided in the Appendix A.3. Figure 3 illustrates that Algorithm 1 generates trajectories with smooth reward curves, closely resembling the learning histories observed in Zisman et al. (2024a). For Task-Level dataset visualization please refer to Appendix D.
@@ 2.2 Cross-Domain Dataset
Table 1 contains detailed statistics of the training dataset. More detailed information on datasets can be found in the Appendix A.
@@
For MuJoCo and Meta-World we utilized demonstrator policies provided in JAT (Gallouédec et al., 2024). However, for several tasks within Meta-World, we trained our own demonstrators due to the unsatisfactory performance of the provided experts. Bi-DexHands demonstrators were trained using official PPO (Schulman et al., 2017) implementation with an increased number of parallel environments. The demonstrator policies for the Industrial Benchmark were trained using the provided scripts, which were built on top of the Stable-Baselines 3 library (Raffin et al., 2021). In-depth view of demonstrators performance can be found in the Appendix G.
@@ 2.3.1 Model architecture
All input sequences are split into groups based on observation and action space dimensionalities. For each group, a separate encoder and decoder MLP head are created, enabling the model to map variable observation and action spaces into a shared embedding space. It is important to note that the model is task-agnostic in the sense that it has access only to the dimensionality-based group identifier, but not to an individual task identifier.
@@ 2.3.2 Training
The input batches are created by collating together multiple input sequences . To standardize the data across tasks, rewards are scaled by task-specific factors (see Appendix F), and observations are normalized. This procedure significantly enhances model performance by reducing task distinguishability based on raw input values, thereby encouraging the model to rely on contextual information to infer the task.
Afterward, sequences are processed by the corresponding encoders to obtain fixed-size representations. The sequence of representations is then passed into a transformer, whose outputs are subsequently decoded into predicted actions . The training objective is to minimize the Mean Squared Error loss between the predicted actions and the ground-truth trajectory actions . Training is conducted on 8 H100 GPUs with a batch size of 64 and 2 gradient accumulation steps. The input sequence length is set to 8192. For more detailed hyperparameter information, refer to Appendix C.
@@ 3.1 Inference-Time Self-Correction on Training Tasks
First, we aim to verify whether the Vintix model has the capability for context-based inference-time adaptation through self-correction. To achieve this, we deploy the model on training tasks (ML45 split for Meta-World, ML20 split for Bi-DexHands and setpoints for Industrial-Benchmark) by iteratively unrolling its actions in a cold-start manner, beginning with an empty initial context. Figure 4 illustrates that the model progressively improves its policy in each domain as the number of shots (episodes played) increases. The agent starts with a suboptimal performance and gradually self-corrects by inferring task-related information from the accumulated context, ultimately reaching near-demonstrator-level performance.
Notably, this improvement is both progressive and consistent along the shots-made axis and is observed across multiple domains with varying dynamics and morphology. For task-level graphs depicting inference-time performance with an empty context, refer to Appendix E.
@@ 3.2 Comparison to Related Action Models
To verify this, we compare the average demonstrator-normalized performance of Vintix with that of Gallouédec et al. (2024) across training tasks in overlapping domains (MuJoCo and Meta-World). JAT scores are computed by aggregating demonstrator-normalized returns over 100 trials per task, as originally reported by the authors. For Vintix, we extract the average performance over 100 episodes after reaching inference-time convergence (i.e., achieving the best -shot performance per task).
Since most of our demonstrators were sourced from JAT, comparing normalized scores is valid. However, we improved expert performance for several tasks by fine-tuning hyper-parameters and normalized our raw returns with respect to these newly enhanced scores. This adjustment lowers our normalized performance relative to JAT, but we adopted this comparison approach as our initial focus is on evaluating the agent’s ability to match demonstrator performance rather than comparing absolute scores.
Figure 5 illustrates that inference-time self-correction enables Vintix to outperform JAT on Meta-World and MuJoCo by significant margins: +32.1% and +13.5%, respectively. These results may further highlight a fundamental advantage of Algorithm Distillation over Expert Distillation, namely its adaptability, as initially reported in Laskin et al. (2022).
@@ 3.3 Generalization Analysis
To evaluate this, we performed the cold-start inference procedure described in previous sections on a set of MuJoCo environments with unobservable variations in viscosity and gravity, which were also unseen during training (for more details on MuJoCo parameter variations, refer to Section A.2.1). Additionally, we applied the same procedure to unseen environments in the Industrial-Benchmark domain (setpoint , see Section A.2.4).
Figure 6 presents the experimental results. As shown, Vintix is still able to achieve near-demonstrator performance in modified environments across both domains. However, a slight decline in convergence quality is observed—slower convergence speed in MuJoCo and subtly diminished asymptotic performance in Industrial-Benchmark. The slower convergence rate suggests that the model requires more iterations of self-correction to reach demonstrator-level policy in environments with moderate parametric variations.
@@
As in the previous experiments, we unroll actions from Vintix with no initial context on test tasks from the Meta-World ML45 split and the Bi-DexHands ML20 split (Section A.2). Overall, we observed that Vintix is not yet capable to handle significantly new tasks. Figure 7 presents one successful rollout and one failure case for each domain (for detailed task-level results, refer to Appendix E). On the Door-Unlock task from Meta-World, Vintix achieved 47% of the mean expert-normalized score, while on the Door-Open-Inward task from Bi-DexHands, it consistently maintained 31% of demonstrator performance. In failure scenarios, the model exhibited random-level performance on the Bin-Picking task and performed below the random baseline on the Hand-Kettle task.
@@ B.1 Training
Appendix C Hyperparameters
@@
Appendix D Task-Level Dataset Visualization
Appendix E Inference Time Performance Graphs
Appendix F Dataset Size and Metadata
| Task | Trajectory Length | Number of trajectories | Mean Episode Length | State space shape | Action space shape | Reward Scaling |
| assembly | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| basketball | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| bin-picking | 100000 | 5 | 100 | (39,) | (4,) | 1.0 |
| box-close | 100000 | 5 | 100 | (39,) | (4,) | 1.0 |
| button-press | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| button-press-topdown | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| button-press-topdown-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| button-press-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| coffee-button | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| coffee-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| coffee-push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| dial-turn | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| disassemble | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| door-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| door-lock | 100000 | 5 | 100 | (39,) | (4,) | 1.0 |
| door-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| door-unlock | 100000 | 5 | 100 | (39,) | (4,) | 1.0 |
| drawer-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| drawer-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| faucet-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| faucet-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| hammer | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| hand-insert | 100000 | 5 | 100 | (39,) | (4,) | 1.0 |
| handle-press | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| handle-press-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| handle-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| handle-pull-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| lever-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| peg-insert-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| peg-unplug-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| pick-out-of-hole | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| pick-place | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| pick-place-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| plate-slide | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| plate-slide-back | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| plate-slide-back-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| plate-slide-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| push-back | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| push-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| reach | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| reach-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| shelf-place | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| soccer | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| stick-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| stick-push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| sweep | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| sweep-into | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |
| window-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 |

### v2

@@ 1 Introduction
In this work, we start an investigation into cross-domain action models through In-Context Reinforcement Learning (ICRL), grounded in the Algorithm Distillation framework (Laskin et al., 2022). By focusing on a data-centric approach to ICRL (summarized in Figure˜2), we present Vintix, a multi-task action model that exhibits initial signs of inference-time self-correction on training tasks (Figure˜1) and preliminary evidence of adaptation to parametric task variations (Section˜3.3).
@@
Continuous Noise Distillation (Section˜2.1): We propose a data collection strategy extending the work of Zisman et al. (2024a) to continuous action spaces to ease the collection of training data.
Open Tools and Datasets for ICRL (Section˜2.2): We publicly release datasets for 87 tasks across four domains (Meta-World, MuJoCo, Bi-DexHands, Industrial-Benchmark), along with data collection tools and instrumentation to support the development of action models eliciting ICRL behavior.
@@ 2 Approach
At the core of our approach (Figure˜2) is Algorithm Distillation (Laskin et al., 2022), a two-step Offline Meta-RL algorithm. The first step involves collecting ordered training histories from base reinforcement learning (RL) algorithms, while the second step involves a decoder-only transformer trained solely for the next-action prediction. This approach facilitates in-context learning by effectively distilling the policy improvement operator into a causal sequence model. We further propose two augmentations to this technique: (1) democratizing the data collection process by introducing a continuous extension of the noise-distillation procedure by Zisman et al. (2024a); (2) conducting generalist agent-style cross-domain training on the acquired dataset.
@@ 2.1 Continuous Noise Distillation
The data collection strategy was originally designed for discrete action spaces. In our work, we propose its extension to continuous action spaces, where the resulting action is defined as a linear mixture of uniform random noise and demonstrator actions. This contrasts with Brown et al. (2019), who employed an epsilon-greedy approach that alternates between fully random and fully expert actions with probability . The pseudo-code for our noise-distillation procedure is formalized in Algorithm˜1.
The noise schedule plays a crucial role in the generation of training trajectories. We observed that linear annealing of often results in non-smooth trajectories with abrupt transitions from random to demonstrator performance. This effect can negatively impact the model’s convergence; therefore, careful tuning of the noise schedule was necessary for certain tasks. A more detailed description of the epsilon decay functions is provided in the Appendix A.3. Figure 3 illustrates that Algorithm 1 generates trajectories with smooth reward curves, closely resembling the learning histories observed in Zisman et al. (2024a). For Task-Level dataset visualization please refer to Appendix E.
@@ 2.2 Cross-Domain Dataset
Table˜1 contains detailed statistics of the training dataset. More detailed information on datasets can be found in the Appendix A.
@@
For MuJoCo and Meta-World we utilized demonstrator policies provided in JAT (Gallouédec et al., 2024). However, for several tasks within Meta-World, we trained our own demonstrators due to the unsatisfactory performance of the provided experts. Bi-DexHands demonstrators were trained using official PPO (Schulman et al., 2017) implementation with an increased number of parallel environments. The demonstrator policies for the Industrial Benchmark were trained using the provided scripts, which were built on top of the Stable-Baselines 3 library (Raffin et al., 2021). In-depth view of demonstrators performance can be found in the Appendix H.
@@ 2.3.1 Model architecture
All input sequences are split into groups based on observation and action space dimensionalities. For each group, a separate encoder and decoder MLP head are created, enabling the model to map variable observation and action spaces into a shared embedding space. It is important to note that the model is task-agnostic in the sense that it has access only to the dimensionality-based group identifier, but not to an individual task identifier. While the task identifier is a unique ID assigned to each task in the dataset, the group identifier simply indicates whether a set of tasks shares the same observation and action space dimensions, as well as the semantic meaning of each channel.
@@ 2.3.2 Training
The input batches are created by collating together multiple input sequences . To standardize the data across tasks, rewards are scaled by task-specific factors (see Appendix G), and observations are normalized. This procedure significantly enhances model performance by reducing task distinguishability based on raw input values, thereby encouraging the model to rely on contextual information to infer the task.
Afterward, sequences are processed by the corresponding encoders to obtain fixed-size representations. The sequence of representations is then passed into a transformer, whose outputs are subsequently decoded into predicted actions . The training objective is to minimize the Mean Squared Error loss between the predicted actions and the ground-truth trajectory actions . Training is conducted on 8 H100 GPUs with a batch size of 64 and 2 gradient accumulation steps. The input sequence length is set to 8192. For more detailed hyperparameter information, refer to Appendix D.
@@ 3.1 Inference-Time Self-Correction on Training Tasks
First, we aim to verify whether the Vintix model has the capability for context-based inference-time adaptation through self-correction. To achieve this, we deploy the model on training tasks (ML45 split for Meta-World, ML20 split for Bi-DexHands and setpoints for Industrial-Benchmark) by iteratively unrolling its actions in a cold-start manner, beginning with an empty initial context. Figure˜4 illustrates that the model progressively improves its policy in each domain as the number of shots (episodes played) increases. The agent starts with a suboptimal performance and gradually self-corrects by inferring task-related information from the accumulated context, ultimately reaching near-demonstrator-level performance.
Notably, this improvement is both progressive and consistent along the shots-made axis and is observed across multiple domains with varying dynamics and morphology. For task-level graphs depicting inference-time performance with an empty context, refer to Appendix˜F.
@@ 3.2 Comparison to Related Action Models
To verify this, we compare the average demonstrator-normalized performance of Vintix with that of Gallouédec et al. (2024) and Sridhar et al. (2024) across training tasks in overlapping domains (MuJoCo and Meta-World). JAT and REGENT scores are taken directly from the respective papers. For Vintix, we extract the average performance over 100 episodes after reaching inference-time convergence (i.e., achieving the best -shot performance per task).
Since most of our demonstrators were sourced from JAT - and REGENT also adopted random and expert scores from JAT - the comparison of normalized scores is valid. However, we improved expert performance for several tasks by fine-tuning hyperparameters and normalized our raw returns with respect to these newly enhanced scores. This adjustment lowers our normalized performance relative to JAT and REGENT, but we adopted this comparison approach as our initial focus is on evaluating the agent’s ability to match demonstrator performance rather than comparing absolute scores.
Figure˜5 illustrates that inference-time self-correction enables Vintix to outperform JAT and REGENT on Meta-World and MuJoCo by significant margins. More detailed comparison can be found in Appendix˜I. These results may further highlight a fundamental advantage of Algorithm Distillation over Expert Distillation, namely its adaptability, as initially reported in Laskin et al. (2022).
@@ 3.3 Generalization Analysis
To evaluate this, we performed the cold-start inference procedure described in previous sections on a set of MuJoCo environments with unobservable variations in viscosity and gravity, which were also unseen during training (for more details on MuJoCo parameter variations, refer to Section˜A.2.1). Additionally, we applied the same procedure to unseen environments in the Industrial-Benchmark domain (setpoint , see Section˜A.2.4).
Figure˜6 presents the experimental results. As shown, Vintix is still able to achieve near-demonstrator performance in modified environments across both domains. However, a slight decline in convergence quality is observed—slower convergence speed in MuJoCo and subtly diminished asymptotic performance in Industrial-Benchmark. The slower convergence rate suggests that the model requires more iterations of self-correction to reach demonstrator-level policy in environments with moderate parametric variations.
@@
As in the previous experiments, we unroll actions from Vintix with no initial context on test tasks from the Meta-World ML45 split and the Bi-DexHands ML20 split (Section˜A.2). Overall, we observed that Vintix is not yet capable to handle significantly new tasks. Figure˜7 presents one successful rollout and one failure case for each domain (for detailed task-level results, refer to Appendix˜F). On the Door-Unlock task from Meta-World, Vintix achieved 47% of the mean expert-normalized score, while on the Door-Open-Inward task from Bi-DexHands, it consistently maintained 31% of demonstrator performance. In failure scenarios, the model exhibited random-level performance on the Bin-Picking task and performed below the random baseline on the Hand-Kettle task.
@@ 4 Related Work
Sequence Modeling in RL.
With the increasing use of Transformers for modeling sequential data, several concurrent works (Chen et al., 2021; Janner et al., 2021) reformulated the Markov Decision Process as a causal sequence modeling problem. Chen et al. (2021) focused on reward conditioning and treated each component of the MDP as a separate token, while Janner et al. (2021) applied beam search over discretized states, actions, and rewards.
Subsequent research has expanded this direction by encouraging such models to maximize returns (Zhuang et al., 2024), adapting Decision Transformers to online learning settings (Zheng et al., 2022), and replacing the Transformer backbone with architectures that scale more efficiently with input length, such as Mamba (Huang et al., 2024).
Another line of work has aimed to scale this modeling paradigm to multi-domain and multi-modal environments (Reed et al., 2022; Gallouédec et al., 2024), or to leverage the in-context learning capabilities of Transformers for Meta-RL (Laskin et al., 2022; Lee et al., 2023). This latter direction is most closely related to Vintix, which is a memory-based, cross-domain Meta-RL method.
@@ B.1 Training
Appendix C Additional Experiments
C.1 Is Algorithm Distillation More Effective Than Expert Distillation?
To assess the benefits of AD-style training on noise-distilled trajectories over simple Expert Distillation (ED), we conducted an experiment using a model with the same architecture as Vintix—including the transformer backbone, encoders, and loss function—but trained exclusively on expert-level demonstrations. We then evaluated both the ED and AD models using the cold-start inference procedure.
Figure˜8 shows that ED underperforms relative to AD in terms of asymptotic performance across both the MuJoCo and Meta-World domains. Specifically, ED converges to an average expert-normalized score of 0.8, whereas AD reaches 0.97 on MuJoCo and 0.95 on Meta-World.
These results highlight the critical role of policy-improvement structure in the dataset for enabling self-correcting behavior and achieving superior performance. Notably, while ED’s performance remains relatively stable across different shot counts in MuJoCo, it shows a positive trend in Meta-World as the number of shots increases.
C.2 Does Vintix Performs In-Context Reinforcement Learning?
To investigate whether Vintix engages in reinforcement learning during inference, we trained a variant of the model without access to reward signals. We then compared its performance to the original Vintix model (trained with rewards) using the cold-start inference procedure on training tasks from both domains.
Evaluation results (Figure˜8) demonstrate that removing the reward signal significantly impairs performance. The reward-masked AD model performs worse asymptotically and exhibits slower convergence on the Meta-World domain.
These findings suggest that reward feedback is essential for effective self-improvement during inference. This supports the hypothesis that supervised training on a dataset containing policy improvement mechanisms enhances the model’s in-context reinforcement learning capabilities.
Appendix D Hyperparameters
@@
Appendix E Task-Level Dataset Visualization
Appendix F Inference Time Performance Graphs
Appendix G Dataset Size and Metadata
| assembly | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| basketball | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| bin-picking | 100000 | 5 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| box-close | 100000 | 5 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| button-press | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| button-press-topdown | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| button-press-topdown-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| button-press-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| coffee-button | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| coffee-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| coffee-push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| dial-turn | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| disassemble | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| door-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| door-lock | 100000 | 5 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| door-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| door-unlock | 100000 | 5 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| drawer-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| drawer-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| faucet-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| faucet-open | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| hammer | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| hand-insert | 100000 | 5 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| handle-press | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| handle-press-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| handle-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| handle-pull-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| lever-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| peg-insert-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| peg-unplug-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| pick-out-of-hole | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| pick-place | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| pick-place-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| plate-slide | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| plate-slide-back | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| plate-slide-back-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| plate-slide-side | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| push-back | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| push-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| reach | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| reach-wall | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| shelf-place | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| soccer | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| stick-pull | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| stick-push | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| sweep | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| sweep-into | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||
| window-close | 100000 | 15 | 100 | (39,) | (4,) | 1.0 | |||||||||||||